You have $269,204 to invest in a stock portfolio (this amount is your original wealth). Your choices are Stock H, with an expected return of 16.3 percent, and Stock L, with an expected return of 10.33 percent. Legal constraints require you to invest at least $49,447 in stock L. If your goal is to create a portfolio with an expected return of 18.58 percent on your original wealth, what is the minimum amount you must borrow (and subsequently repay) at the risk free rate of 2.31 percent to achieve your goal?
Answer #1

Let the amount borrowed be x

Portfolio return is equal to weighted average return

49447*10.33% + (269204+x-49447)*16.3% - x*2.31% = 269204*18.58%

5107.8751 + 35,820.391 + 0.1630x – 0.0231x = 50,018.1032

0.1399x = 9089.8371

X = $64,973.82

Hence, amount required to be borrowed = $64,973.82